OpenStudy (anonymous):

Merp help please <3 Which of the following is a factor of x2 - 5x - 24? x + 3 x - 6 x + 4 None of the above

OpenStudy (anonymous):

x² - 2x - 24 = 0 (x - 6)(x + 4) = 0 (x - 6) and (x + 4) are the two factors x = 6 or x = -4 are the two roots im so confused did i do this wrong?

OpenStudy (anonymous):

that second one was done perfectly well

OpenStudy (anonymous):

seccond one? oh! i mixed up the 5 with a 2 no wonder!

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

was the original written wrong?

OpenStudy (anonymous):

or the second one?

OpenStudy (agent0smith):

Just find two numbers that multiply to -24 and add up to 5... -8 and +3. Also, if x+3 is a factor, then x=-3 is a zero if x-6 is a factor, then x=6 is a zero if x+4 is a factor, then x=-4 is a zero. So if you plug in those x values, and the x2 - 5x - 24 equals zero, then you've found your factor